# Stylometry recognizes human and LLM-generated texts in short samples

### v1

@@ 4.4.1 Multiclass classification
SHAP explanations were collected and averaged over all cross-validation folds. In Fig. 1 the ten most important StyloMetrix and frequency features are shown.
The StyloMetrix features include (in the order of importance): number of function word types, number of words in narrative sentences, the type-token ratio for words lemmas, statistics between noun phrases, fronting, difference between the number of words and the number of sentences, punctuation – dots, punctuation, punctuation – commas, and numerals; see (Okulska et al., 2023) for feature descriptions. The frequency features include single part-of-speech tags such as: whitespace, nouns, adpositions, proper nouns, verbs, adjectives, and determiners; POS bigrams such as: noun followed by a whitespace; and single lemmas such as: ’despite’, ’and’.
@@
Figure 2: Local explanations of 10 most important StyloMetrix (a-c) and frequency features (d-f) in multiclass classification for text samples describing the term ‘The Swarbriggs’. Only selected models are shown. For this term, the Wikipedia was classified correctly, GPT-4 was misclassified as the Wikipedia, and Llama2 was misclassified as Orca. Grey numbers to the left indicate feature value in this particular text sample. The positive/negative SHAP values do not point strictly to any particular class (in the multiclass scenario) but they tend to be higher for Wikipedia and GPT models and lower for worse models.
@@
Notice the dates in the Wikipedia sample (POS_NUM), lower number of punctuation marks for Llama2 than for the Wikipedia (see numbers next to L_PUNCT in Fig. 2a), SENT_D_NP having similar values in all three cases. Also, looking at Fig. 2b, one notices a significantly larger number of proper nouns and dates in Wikipedia (PROPN – also in bigrams – and NUM), redundant spaces in Llama2 (SPACE), and other singular features. It is worth recalling that models trained on different data subsets (CV folds) contribute to the SHAP values in Fig. 1, while the SHAP values presented in Fig. 2) correspond to a single classifier whose test set contained the selected texts. Also, the SHAP values in Fig. 2) are averaged over classes, but one can obtain explanations for each class separately.
Text samples (corresponding to the term ‘The Swarbriggs’) are shown in Fig. 3-5, where also the frequency features most important to the classifier have been marked.
4.4.2 Binary classification
Here we present only the example of classifying the Wikipedia and GPT-4, as shown in Figures 6, respectively, for StyloMetrix and frequency features. Analogous analyses can be repeated for the other pairs of classes. Let us recall, that punctuation (including the SPACE token) was excluded from the frequency features. Like above in the multiclass scenario, one notices features representing proper names (L_PROPER_NAME, PROPN), dates and other numerals (POS_NUM, NUM), etc. GPT-4 strikingly tends to abuse words like ‘significant’, ‘notable’ or ‘despite’. Its usage of grammatical features (i.e., POS n-grams), however, tends to be strongly frequency-standardised, visible as the red bulks of the distributions in contrast to the long grey outlying distributions for the Wikipedia.
Figure 6: Explanations for binary classification between the Wikipedia and GPT-4 using StyloMetrix (a) and frequency-based (b) features. Only the first 10 most important features are shown. Each point is a 10-sentence sample describing a given term coloured by: (Left) the sample’s class, and (Right) its feature’s intensity. Positive SHAPs point toward GPT and negative ones toward the real Wikipedia.
@@ 4.5 Summarization methods comparison
The text summarization methods are used only for comparison reasons what popular methods perform in a 1vs1 classification against language models. Similar to LGBM experiments, this one was also performed on the first prompt, because of not significant differences between both prompts in the decision tree classification. The classification of summarization methods was also performed using decision tree method. The results are given in Table 5.
Wikipedia sum sumy ts bart
@@
gpt3.5 0.7540 0.8398 0.8889 0.9648
gpt4 0.7283 0.8071 0.8967 0.9501
llama2 0.8924 0.9129 0.9034 0.8135
llama3 0.6865 0.79 0.8757 0.9622
orca 0.9046 0.9223 0.9107 0.7561
falcon 0.8362 0.8875 0.8353 0.7935
@@
The worsts recognized model is GPT-4 as the comparison with the Wikipedia summary is only on about 72%. This indicates that this model can simulate the way how human summarize the Wikipedia pages, but it is important to highlight that is was also the most complex model used in our experiment. The other questionable recognitions were obtained for Orca vs. BART summarizer and Sumy summarizer vs. Wikipedia, about 75% both. The other results vary between 80% and 92%. The best results were achieved by 1vs comparisons as follows: T5 summarizer vs. Sumy summarizer – about 92%, Sumy summarizer vs. Orca – about 92%, GPT-4 vs. BART summarizer – about 95%, Llama3 vs. BART summarizer – about 96%, BART summarizer vs. Wikipedia – about 96%, and BART summarizer vs. Sumy summarizer – about 97%.
@@ 5 Discussion
From multiclass explanations: it seems that good models do not have single strongly recognisable features, but their style is more dispersed among the quantified features. Interestingly, simple features like the number of punctuation marks matter. The whitespaces found in Llama2 were actually double spaces between tokens or a space at the beginning of the text. The number of full stops appears as a distinguishing feature possibly because the LLMs tend to stop generating the text in the middle of the sentence. This might also affect ‘the difference between the number of words and the number of sentences’ (SENT_ST_WRDSPERSENT) as well as some other features. Wikipedia descriptions tend to be more fact-packed (dates and proper nouns) than LLM-generated ones. The distributional plots from binary classification between Wikipedia and GPT-4, suggest that the LLM favours certain individual words and is more standardised than Wikipedia in terms of grammatical structures (represented by frequencies of part-of-speech n-grams) – perhaps an expected outcome since the Wikipedia text sample were authored by many people.
The summarization methods achieve similar results as for the decision tree experiment. We can conclude that we will achieve similar results as for LGBM for the summarization methods. It indicates that summarization methods does have distinctive way of text summarization that can be found using stylometry.
@@ 5.1 Limitations
The limitations of the present paper concern mainly the material of the analysis. Firstly, the results and specific conclusions refer only to the chosen text type, i.e., introductions to Wikipedia articles, which are expected to conform to an encyclopaedic style: plain, factual and partly formulaic. Some of the most distinctive features reflect that, and cannot be generalised to classifying other text types. However, the analytic pipeline is generic, including the engineered features, which have been designed and used in the context of literary texts.
Secondly, the language of the text samples is limited to English only. The precise lexical, grammatical and other complex features will differ for other languages. Performance of stylometric tools has been known to depend heavily on language and specifically on language type (analytic, synthetic, etc.), see, e.g. (Eder, 2011, Evert et al., 2017). However, the LLMs are also best developed in English (Li et al., 2024) and hence we expect it to be the most challenging setting for classification. The text processing pipeline we used strictly depends on the availability of NLP tools (like POS taggers, dependency parsers, NERs, etc.) for a given language. The frequency features at this moment depend on spaCy, which currently provides more or fewer tools for about 24 languages. In the case of StyloMetrix features, even though they also depend on the models distributed by spaCy, they were custom-designed for Polish, English, German, Ukrainian and Russian only.
@@ 6 Further works
The results show we can use stylometry for english language to distinguish between large language models and human written text. The next steps would be to perform the analysis on different languages, including languages used by a rather small number of people in total.
The second way of extending this research is to use other stylometry libraries, classification methods, and more complex language models. Based on the presented results, the more complex models shows that they are harder to be differentiate from human written text compared to the less complex models.
The third vector of further research is to extend the features list and add other features like fractal based features. As stylometry seems to be a good choice, there might be other ones that might be more precise.
@@
The notebook with the source code can be found in the repository:
@@
It includes the URLs to the libraries we used, the Python code to get the data, preprocess it, and execute the experiment. It comes with a enviornment setup guidelines.
@@ Acknowledgements
The research for this publication has been supported by a grant from the Priority Research Area DigiWorld under the Strategic Programme Excellence Initiative at Jagiellonian University. JKO’s research on the stylometric pipeline was financed by the European Regional Development Fund as a part of the 2014–2020 Smart Growth Operational Programme, CLARIN—Common Language Resources and Technology Infrastructure, project no. POIR.04.02.00-00C002/19.

### v2

@@ 4.5 Explainability
This section reports the results of SHAP explanations. In the binary classification, we provide only a single example to show the effectiveness of the explanations. For that purpose the GPT-4 model was chosen as the hardest one to detect. In this case, the positive or negative direction of SHAP values points toward one or the other class. In the multiclass classification, the obtained explanations take into account all the LLMs. In that case, the absolute values of SHAP show which features explain which model to what extent. Depending on the model’s idiosyncrasies, a feature can explain several models well, but others poorly. Moreover, some models may be explained by few very strong features (i.e., with large SHAP values), while others may need numerous features contributing only small fractions to the explanation, as visible in Figure 2.
For each classification scenario, SHAPs were collected and averaged across all CV folds.
4.5.1 Binary classification
Here we present only the example of classifying the Wikipedia and GPT-4, as shown in Figures 1(a) and (b), respectively, for StyloMetrix and frequency features. Analogous analyses can be repeated for the other pairs of classes. Let us recall, that punctuation (including the SPACE token) was excluded from the frequency features. Like above in the multiclass scenario, one notices features representing proper names (L_PROPER_NAME, PROPN), dates and other numerals (POS_NUM, NUM), etc. GPT-4 strikingly tends to abuse words like ‘significant’, ‘notable’ or ‘despite’. Its usage of grammatical features (i.e., POS n-grams), however, tends to be strongly frequency-standardised, visible as the red bulks of the distributions in contrast to the long grey outlying distributions for the Wikipedia.
Figure 1: Explanations for binary classification between the Wikipedia and GPT-4 using (a) StyloMetrix and (b) frequency-based features. Only the first 10 most important features are shown. Each point is a 10-sentence sample describing a given term coloured by: (Left) the sample’s class, and (Right) its feature’s intensity. The left plots indicate whether positive or negative SHAPs point toward GPT or the real Wikipedia.
@@ 4.5.2 Multiclass classification
In Fig. 2 the ten most important StyloMetrix and frequency features are shown.
The StyloMetrix features include (in the order of importance): number of function word types, number of words in narrative sentences, the type-token ratio for word lemmas, statistics between noun phrases, fronting, difference between the number of words and the number of sentences, punctuation – dots, punctuation, punctuation – commas, and numerals; see (Okulska et al., 2023) for feature descriptions. The frequency features include single part-of-speech tags such as: whitespace, nouns, adpositions, proper nouns, verbs, adjectives, and determiners; POS bigrams such as: noun followed by a whitespace; and single lemmas such as: ‘despite’, ‘and’.
@@
Figure 3: Local explanations of 10 most important StyloMetrix (a-c) and frequency features (d-f) in multiclass classification for text samples describing the term ‘The Swarbriggs’. Only selected models are shown. For this term, the Wikipedia was classified correctly, GPT-4 was misclassified as the Wikipedia, and LLaMa 2 was misclassified as Orca. Grey numbers to the left indicate feature values in this particular text sample. The positive/negative SHAP values do not point strictly to any particular class (in the multiclass scenario) but they tend to be higher for Wikipedia and GPT models and lower for worse models.
@@
The feature explanations are not model-agnostic. Notice the dates in the Wikipedia sample (POS_NUM), lower number of punctuation marks for LLaMa 2 than for Wikipedia (see numbers next to L_PUNCT in Fig. 3a), SENT_D_NP having similar values in all three cases. Also, looking at Fig. 3b, one notices a significantly larger number of proper nouns and dates in Wikipedia (PROPN – also in bigrams – and NUM), redundant spaces in LLaMa 2 (SPACE), and other singular features. One can notice that the explanations for the GPT models are more distributed (no single feature with a huge SHAP value) than the other models.
It is worth recalling that models trained on different data subsets (CV folds) contribute to the SHAP values in Fig. 2, while the SHAP values presented in Fig. 3 correspond to a single classifier whose test set contained the selected texts. In addition, the SHAP values in Fig. 3 are averaged over classes, but explanations for each class can be obtained separately.
Text samples (corresponding to the term ‘The Swarbriggs’) are shown in Fig. 4-6, where also the frequency features most important to the classifier have been marked.
@@ 4.6 Summarization methods comparison
The text summarization methods are used only for comparison reasons as to what popular methods perform in a binary classification against language models. Similarly to LGBM experiments, this was also performed on the first prompt, because there were no significant differences between both prompts in the decision tree classification. The classification of summarization methods was also performed using the decision tree method. The results are given in Table 10.
Wikipedia sum Sumy T5 BART Gensim
@@
Gensim 0.9611 0.9753 0.948 0.7954 1.0
GPT-3.5 0.7540 0.8398 0.8889 0.9648 0.9634
GPT-4 0.7283 0.8071 0.8967 0.9501 0.9469
LLaMa 2 0.8924 0.9129 0.9034 0.8135 0.8986
LLaMa 3 0.6865 0.79 0.8757 0.9622 0.9509
Orca 0.9046 0.9223 0.9107 0.7561 0.8717
Falcon 0.8362 0.8875 0.8353 0.7935 0.8769
@@
The worst-recognized model is GPT-4, as the comparison with the Wikipedia summary is only about 72%. This indicates that this model can simulate the way human summarizes Wikipedia pages, but it is important to highlight that it was also the most complex model used in our experiment. The other questionable recognitions were obtained for Orca compared to BART summarizer and Sumy summarizer compared to Wikipedia, about 75% both. The other results vary between 80% and 92%. The best results were achieved in a binary classification to recognize the text generated by one summarizing methods from another summarizing method or LLM, like T5 summarizer and Sumy summarizer – with about 92%, Sumy summarizer and Orca – about 92%, GPT-4 and BART summarizer – about 95%, LLaMa 3 and BART summarizer – about 96%, BART summarizer and Wikipedia – about 96%, and BART summarizer and Sumy summarizer – about 97%.
@@ 5 Discussion
From multiclass explanations: it seems that well-performing models do not have single strongly recognisable features, but their style is more dispersed among many quantified features. Moreover, the explanations are not general, but may vary depending on the model, hence, the multiclass training is indispensable. These plots summarise all folds in the cross-validation loop, so the results are also stable in terms of different training/test splits. Interestingly, simple features such as the number of punctuation marks matter. The whitespaces found in LLaMa 2 were actually double spaces between tokens or a space at the beginning of the text. The number of full stops appears as a distinguishing feature, possibly because the LLMs tend to stop generating the text in the middle of the sentence. This might also affect ‘the difference between the number of words and the number of sentences’ (SENT_ST_WRDSPERSENT) as well as some other features. Wikipedia descriptions tend to be more fact-packed (dates and proper nouns) than LLM-generated ones. The distributional plots from binary classification between Wikipedia and GPT-4, suggest that the LLM favours certain individual words and is more standardised than Wikipedia in terms of grammatical structures (represented by frequencies of part-of-speech n-grams) – perhaps an expected outcome since the Wikipedia text samples were authored by many people. These conclusions come from explanations collected in the cross-validation loop, so they are stable in terms of different training/test splits.
The summarisation methods achieve similar results in the decision tree experiment. We can conclude that we will achieve similar results in LGBM for the summarisation methods. It indicates that the summarisers do have a distinctive way of text summarisation that can be found using stylometry. Succinctness of BART and artefacts in T5 explains their high recognisability. Sumy is the most successful due to its flexibility in choosing the summary length.
@@ 5.1 Limitations
The limitations of the present paper concern mainly the material of the analysis. Firstly, the results and specific conclusions refer only to the chosen text type, i.e., introductions to Wikipedia articles, which are expected to conform to an encyclopaedic style: plain, factual and partly formulaic. Some of the most distinctive features reflect that, and cannot be generalised to classifying other text types. However, the analytic pipeline is generic, including the engineered features, which have been designed and used in the context of literary texts. Whether the cross-domain classification with this type of model is robust is at this time debatable, taking into account our preliminary results in the AuTexTification task Sarvazyan et al., 2023a, but also results of others that have tried utilising stylometric features Mikros et al., 2023 with results below simple baselines. One can frame the issue of domain dependence of the model in various ways: both training and testing in another domain or cross-domain detection (i.e., detection on unseen domains), and which part of the data is unseen (whether the human- or machine-generated texts or both). Depending on these choices, the attack scenario is more or less realistic.
Secondly, the language of the text samples is limited to English only. The precise lexical, grammatical and other complex features will differ for other languages. Performance of stylometric tools has been known to depend heavily on language and specifically on language type (analytic, synthetic, etc.), see, e.g. (Eder, 2011, Evert et al., 2017). However, the LLMs are also best developed in English (Li et al., 2024b) and hence we expect it to be the most challenging setting for classification. The text processing pipeline we used strictly depends on the availability of NLP tools (like POS taggers, dependency parsers, NERs, etc.) for a given language. The frequency features at this moment depend on spaCy, which currently provides more or fewer tools for about 24 languages. In the case of StyloMetrix features, even though they also depend on the models distributed by spaCy, they were custom-designed for Polish, English, German, Ukrainian and Russian only.
@@
Besides the single domain constraint, the robustness testing of the stylometric detection methods and their explanations is limited in terms of the variability of LLM generation. One can envisage generating multiple text versions with: varying hand-crafted or machine-paraphrased prompts, persona-assigned prompts Przystalski et al. (2025), Liu et al. (2024a), Wang et al. (2024a), as well as the same prompt with varying LLM parameters. In our Wikipedia-based dataset, however, we do not expect much variance by varying the prompts, due to the constraints of the encyclopaedic style. In this case, we consider varying and tuning the prompts a less realistic attack scenario.
One should also note, that the multiclass classification is performed on a closed set of classes. Although adding unseen models does not change the task in binary classification (human vs. machine), in the multiclass case the task would change to an open set problem.
While the binary classification task (human vs. machine detection) remains unchanged with the addition of new generative models, the multiclass setting fundamentally changes: the task becomes an open-set classification problem Geng et al. (2020), where the classifier additionally has to recognize samples that belong to unknown or novel classes. This issue is out of scope of the present paper, however, having a good close-set classifer is helpful in the open-set problems Vaze et al. (2022).
@@
In the subtask 1 of “Voight-Kampff Generative AI Detection at PAN and ELOQUENT 2025” (Bevendorff et al., 2025) (essays, news, and fiction genres as well as their obsfuscated versions) our pipeline (Ochab et al., 2025b) without hyperparameter optimisation has reached $F_{1}=0.823$ against the top result $F_{1}=0.898$. A recent zero-shot detection solution achieved in Sun & Lv (2025) accuracy of 90,6%. The average accuracy on three different structured texts datasets is 79,26%. Both show that stylometric approach achieves better results. In Xu et al. (2024) FreqMark method was proposed for LLM generated text using frequency-based watermark. It shows robustness against paraphrasing and other attack methods. The accuracy of 98% shows that not only stylometry-based methods perform well on paraphrased text. Stylometric methods were used in Al-Shaibani & Ahmed (2025) for text fingerprints to detect text generated by LLMs. For four models: two Arabic and two general text, the accuracy vary for LLM models text generation detection between 88.23 and 98.07% for social media content.
Model Falcon GPT-3 GPT-4 LLaMA 2 LLaMA 3 Orca Human
GPTZero prompt #1 98% 100% 96% 98% 98% 98% 100%
GPTZero prompt #2 100% 99% 93% 95% 100% 99% N/A
HIX prompt #1 3% 5% 3% 5% 4% 0% 100%
Table 11: Commercial applications predictions
5.2 Commercial applications
We have tested two different commercial solutions. The test set was a randomly chosen set of 100 prompt results for each model separately. We also included 100 human-written text. The number of chosen samples is caused due to the high costs of each request of such tools. The results are presented in Table 11. noa (2025) performs very well and predicts if the text is written by a human or a model almost perfectly. The disadvantage of this solution, similar to almost every commercial solution, is that it is a binary classification: human or AI. The models correctly identified the generated text as LLM generated, but there are no details on exactly what LLM was used for generation. Our proposed solution performs slightly worse than GPTZero, but it is a multi-label classifier. For comparison, we have tested the classifier developed by HIX A.I. (2025). It performs very poorly and classifies almost every sample as written by humans.
@@ 6 Further works
The results show that we can use stylometry for the English language to distinguish between LLMs and human written text. The next steps would be to perform the analysis in other languages, including low-resource languages.
The second way to extend this research is to use other stylometry libraries, classification methods, and more complex language models. Based on the results presented, the more complex models show that they are harder to differentiate from human written text compared to the less complex models.
The third vector of further research is to extend the feature list with features encoding long memory and correlations in text, such as fractal-based features. As stylometry seems to be a good choice, there might be other ones that might be more precise.
Another extension can be the use of stylometry together with neural embeddings. A hybrid approach might increase the accuracy in generated text recognition.
Finally, the classification explanations obtained by different methods and from different classifiers should be verified for their consistency and stability across various domains.
@@
The data files and code used for text preprocessing and analysis can be found at https://osf.io/dfz6k/ (Ochab et al., 2025a).
The source code for text generation can be found in the repository:
@@
It includes: the URLs to the libraries, the code to get the data, preprocess it, and execute the experiment. It comes with setup guidelines, contains all parameters set for each model.
@@ Acknowledgements
We thank Tomasz Walkowiak for many insightful comments.
The research for this publication has been supported by a grant from the Priority Research Area DigiWorld under the Strategic Programme Excellence Initiative at Jagiellonian University.
JKO’s research on the stylometric pipeline was financed by European Funds for Smart Economy, FENG program, CLARIN – Common Language Resources and Technology Infrastructure, project no.FENG.02.04-IP.040004/24-00.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857533 and from the International Research Agendas Programme of the Foundation for Polish Science No MAB PLUS/2019/13.
The publication was created within the project of the Minister of Science and Higher Education “Support for the activity of Centers of Excellence established in Poland under Horizon 2020” on the basis of the contract number MEiN/2023/DIR/3796.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used GPT and Writefull LLM models to improve the readability, style and grammar of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.